The population x(t) of a certain city satisfies the logistic law dx/dt = (x÷100)- (x^2 ÷ 10^8)
where t is measured in years. Given that the population of the city is 100000 in 1980, determine the population at any time t >1980 . Also find the population in the year 2000.
1
Expert's answer
2020-04-03T09:37:11-0400

"\frac{dx}{dt}=\frac{x}{100}-\frac{x^2}{10^8}"

"10^8\frac{dx}{10^6x-x^2}=dt"

"10^8\int{\frac{dx}{10^6x-x^2}}=\int{dt}"

"100ln\frac{x}{x-10^6}=t+C_1"

"\frac{x}{x-10^6}=Ce^{\frac{t}{100}}"

Given that in the year 1980, t=0 and "x(0)=10^5" we get "C=-\frac{1}{9}"

Then, "\frac{x}{x-10^6}=-\frac{1}{9}e^{\frac{t}{100}}"

"x(t)=\frac{10^6e^{\frac{t}{100}}}{e^{\frac{t}{100}}+9}"

in the year 2000, t=20

"x(20)=\frac{10^6e^{\frac{20}{100}}}{e^{\frac{20}{100}}+9}=119,494.6"